OpenStudy (anonymous):

When a chunk of matal is dropped into a container of water the water level rises 3.4cm. The container is shaped like a cylinder with a radius of 3cm. What is the volume of the metal. Show work please<33

OpenStudy (anonymous):

pi r^2 h

OpenStudy (anonymous):

Volume of cylinder = (Area of the Base)*(Height) The base of a cylinder is a circle so, Area of a circle = (pi)r^2 The initial Height is X, after the metal is dropped it turns to x+3.4

OpenStudy (anonymous):

soo how would you find the volume of that metal? or the inital height?

OpenStudy (anonymous):

it is same as volue of water risen in container

OpenStudy (anonymous):

how would you find that out?

OpenStudy (anonymous):

\[V' = \pi (3^{3})h = 9{\pi}h\]\[V'' = \pi(3^{3})(h+3) = 9h{\pi}+27{\pi}\] \[V'' - V' = Vmetal\]\[9h{\pi}+27{\pi} - 9{\pi}h = 27{\pi}\]

OpenStudy (anonymous):

Where V' = original volume V'' = after metal volume

OpenStudy (anonymous):

oh okay, but on the first equation for the after metal volume shouldnt it be (h+3.4)? or no..

OpenStudy (anonymous):

yea, you're right. instead of 27 it's 30.6

OpenStudy (anonymous):

so its 30.6pi?

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

Thanks so much I've got it now<3

OpenStudy (anonymous):

np